OpenStudy (anonymous):

Two Jupiter-size planets are released from rest 1.40×1011m m apart. What are their speeds as they crash together?

OpenStudy (anonymous):

Hold on is the 1011 10^11 or just 1011???

OpenStudy (anonymous):

10^11

OpenStudy (anonymous):

M₁ = mass of planet #1 M₂ = mass of planet #2 M = total mass R₁ = radius of planet #1 R₂ = radius of planet #2 d₁ = initial distance between planet centers d₂ = final distance between planet centers a = semimajor axis of plunge orbit v₁ = relative speed of approach at distance d₁ v₂ = relative speed of approach at distance d₂ M₁ = M₂ = 1.8986e27 kilograms M = M₁ + M₂ = 3.7972e27 kg G = 6.6742e-11 m³ kg⁻¹ sec⁻² GM = 2.5343e17 m³ sec⁻² d₁ = 1.4e11 meters a = d₁/2 = 7e10 meters R₁ = R₂ = 7.1492e7 meters d₂ = R₁ + R₂ = 1.42984e8 meters v₁ = 0 v₂ = √[GM(2/d₂−1/a)] v₂ = 59508.4 m/s The time to fall is 1337.7 days :)) Does this help of is it just confzzing???? :))

OpenStudy (anonymous):

The top is what the variables mean and the bottom is the math!!! ;))

OpenStudy (anonymous):

Or think of it like this (I have reasoning in this one! F = GMm/r^2 where M = m = mass of Jupiter = 1.8987 × 10^27 kilograms r = 1.40 x 10^11 kilometers = 1.40 x 10^14 meters (so the answer will come out in Newtons) G = 6.67 × 10−11 N · m2/kg^2 mass^2 = 3.598609 x10^54 kg2 r^2 = 1.96 x 10^28 meters ^2 F= ((6.67 × 10−11 N · m2/kg2)(1.8987 × 10^27 kilograms)^2)/(1.40 x 10^14 meters )^2 acceleration = 0 m^2/sec^2, velocity = 0 m/sec,since the planets were released at REST, which means both velocity and acceleration = 0. All acceleration is due to the mass and gravitational acceleration due to the mass of the two planets m^2 = 3.60506169 x 10^54 kg^2 r^2 = 1.96 x 10^28 m^2 F = 1.226824565 x 10^16 Newtons Divide by the mass x G to get the acceleration. If you want the velocity, you have to integrate the acceleration according the basic formula y = x + v + a where x and y = coordinates on a two dimensional plane, v = velocity a = acceleration. First derivative =dy /dx = x^0 + v, second derivative = d^2y/dx^2 = a. This is basic second semester calculus, and physics is applied math lol.

OpenStudy (anonymous):

Do you understand?????

OpenStudy (anonymous):

I understand what you are saying. I've been stuck on this problem for a while and just tried the answer. Unfortunately, it's not right :/

OpenStudy (anonymous):

No its right........ I already had this quesion on one of my previous assensments and I got it correct.... :\

OpenStudy (anonymous):

Hold on what did you put in for your answer?

OpenStudy (anonymous):

59508m/s

OpenStudy (anonymous):

Did you read my answer or just skip to the bottom?????? Because that not correct.....

OpenStudy (anonymous):

Divide by the mass x G to get the acceleration. If you want the velocity, you have to integrate the acceleration according the basic formula y = x + v + a where x and y = coordinates on a two dimensional plane, v = velocity a = acceleration. First derivative =dy /dx = x^0 + v, second derivative = d^2y/dx^2 = a.

OpenStudy (anonymous):

I think you took a pre answer hold on........ well just keep in mind the process... :))

OpenStudy (anonymous):

I read the whole thing and thought I understood it. It made sense for the most part. Your determination of the a value did not make sense to m though. I have been trying to solve it with only using conservation of energy.

OpenStudy (anonymous):

Ok see I solve things differently, and usually it takes longer but it works for me, yeah know , well im glas your read the whole thing! Most people don't and it is soo annoying!! :))

OpenStudy (anonymous):

I'm pretty sure it'll be on my final so I need to make sure that I know it :P

OpenStudy (anonymous):

Oh well good luck! I don't have to take those yet see im in 7th grade, but if I wanted to I could skip ALL of high school!!!!! :)) So happy I just found out this morning! :))

OpenStudy (anonymous):

Thanks :) NowI just need to figure out how to get the right answer. Or at least get it and work backwards

OpenStudy (anonymous):

Ok well I need to go, I have homework, do you want a website that I use all the time for calc help????

OpenStudy (anonymous):

These are sooo great! They explain things soo well!!!!! :)) There also the reason I did not fail calc 1 and 2!

OpenStudy (anonymous):

thanks :)